can you solve this
one angle of a triangle is 8 greater thanthe smallest angle. the third angle is 4 greater than twice the smallest angle. find the measure of the three angles.

Answer by rfer(16322)

x+x+8+2x+4=180
4x=168
x=42
x+8=50
2x+4=88